NBC company is expected to pay a $2.50 per share dividend at the end of this year (i.e., D1=$2.50). The dividend is expected to grow at a constant rate of 6% a year from the current period to the foreseable future. The required rate of return on the stock, rs , is 15%. What is the price per share of NBC's stock today?

1.

$25.00

2.

$27.78

3.

$25.75

4.

$20.00
Answer #1

P0 = D1 / (k - g)

= 2.50 / (0.15 - 0.06)

= 2.50 / 0.09

P0 = $ 27.78

Option B is correct.